In the linear system 2m+5n=14 2m-3n=6, how do I solve? Which variable should I solve for?
Answer by jim_thompson5910(35256)

Hint: Subtract the second equation from the first one to get

2m+5n = 14
2m-3n = 6
----------
0m+8n = 8

So
8n = 8

Solve for n. Then use this to find m.